Calculate the volume of 0.660-M NaOH solution needed to completely neutralize 49.6 mL of a 0.460-M solution of the diprotic acid H2C2O4.
Answer #1

Balanced chemical equation is:

H2C2O4 + 2 NaOH ---> Na2C2O4 + 2 H2O

Here:

M(H2C2O4)=0.46 M

M(NaOH)=0.66 M

V(H2C2O4)=49.6 mL

According to balanced reaction:

2*number of mol of H2C2O4 =1*number of mol of NaOH

2*M(H2C2O4)*V(H2C2O4) =1*M(NaOH)*V(NaOH)

2*0.46 M *49.6 mL = 1*0.66M *V(NaOH)

V(NaOH) = 69.1 mL

Answer: 69.1 mL